cathrinerose:

If f(x) = 8 – 10x and g(x) = 5x + 4, what is the value of (fg)(–2)?

cathrinerose:

can someone help me with this

dude:

Question, is the `(fg)(-2)`  (f o g)?

cathrinerose:

imma try thx

dude:

I am asking you if that is in the question

dude:

I want to be sure if it's a composite function

cathrinerose:

oh the question is if f(x)=8-10x and g(x)=5+4 then what is the value of (fg)(-2)

dude:

No symbol in between then?

cathrinerose:

i dont think so

dude:

Okay so you are multiplying both functions \((8-10x)\times (5x+4)\) Now substitute -2 in for x \((8-10(-2))\times (5(-2)+4)\)

dude:

Do you know how to solve from there?

cathrinerose:

oka so ur just distributing right

dude:

Sure, make sure to solve for what is inside the parenthesis first

cathrinerose:

ok oka thx

dude:

No problem

cathrinerose:

ur a cool guy

dude:

Thank you

cathrinerose:

np

cathrinerose:

thx for the medal

dude:

Of course

cathrinerose:

:)